OpenStudy (anonymous):

A crane has a power of 54 kW. It takes the crane 65 s to lift the large container. How much work is generated by the load? (HINT: Convert from kW to W first)

OpenStudy (shamim):

Given p=54kw=?w t=65sec w=?

OpenStudy (shamim):

Required formula is p=w/t w=?

OpenStudy (shamim):

Response plz!!!!!!!!

OpenStudy (anonymous):

Okay so I am not sure if this is the answer but I did this: 54=w/65 So I divided the 65 by the 54 and I got 0.83

OpenStudy (shamim):

U know p=54kilowatt=54*1000watt=54000W Right??

OpenStudy (shamim):

Because 1kw=1000wat

OpenStudy (shamim):

We can write watt by only w

OpenStudy (shamim):

Plz feel free to ask

OpenStudy (shamim):

Medal?????!!!!!!!!